# BATIS: Bayesian Approaches for Targeted Improvement of Species Distribution Models

### v1

@@ 1 Introduction
Understanding the factors governing the spatial distribution of species is a long-standing topic of interest in ecology (de Candolle 1855; Thuiller et al. 2015). The rapid loss of biodiversity further highlights the need to better understand species distributions and conservation status, which requires robust modeling methods (Sharma et al. 2025). Species distribution models (SDMs) are at the forefront of this research, enabling ecologists to predict and understand species occurrence by linking observational data (presence or absence) to environmental covariates (Elith and Leathwick 2009). These models have proved to be essential tools to support conservation planning and management (Jetz et al. 2019), with notable examples including predicting species range shift (Chauvier-Mendes et al. 2024), mapping invasive species risk (Davis et al. 2024), supporting translocation programs (Draper, Marques, and Iriondo 2019), reducing bycatch rates in fisheries (Hazen et al. 2018), and informing IUCN Red List assessments (Syfert et al. 2014).
SDMs are traditionally based on statistical approaches, offering strong interpretability and theoretical grounding. However these methods often rely on restrictive assumptions such as linearity, leading them to struggle with high-dimensional datasets. Recently, deep learning approaches to SDMs proposed to overcome these challenges and were shown to consistently outperform statistical ecology methods for large and heterogeneous datasets (Deneu et al. 2021).
Although a wide variety of SDM approaches have now been introduced, their effectiveness at large spatio-temporal scales remains hindered by the abundance and geographical coverage of the data used to train them (Van Proosdij et al. 2016). In most cases, observations are disproportionately concentrated in areas easily accessible by roads (Kadmon, Farber, and Danin 2004) and with high human population densities (Speed et al. 2018). In contrast, remote locations are most often associated with very limited to no observations, as illustrated in Fig. 1.
@@
Data distribution biases especially affect machine learning-based approaches, which typically rely on trustworthy labeled data. Limited observations in a given location can lead to poor estimates of species occurrence (Fink et al. 2020) – essentially, noisy labels – and naïvely incorporating such data points in model training can degrade performance. However, sparse observations still offer potentially valuable information, and finding ways to properly leverage it could help improve SDMs and guide conservation efforts in remote or underresourced areas. We argue that uncertainty-aware methods can offer a principled way of leveraging limited location-specific observations, enabling controlled prediction updates that can balance global-scale models with local-scale observations.
@@
  • We benchmark a variety of state-of-the-art uncertainty estimation ML methods on the above framework, and present a novel dataset derived from eBird (Sullivan et al. 2009), building on the SatBird (Teng et al. 2023) dataset.
@@ 2 Related Works
Species distributions have been widely estimated from environmental data with statistical ecology methods. Popular SDMs in ecology include techniques based on maximum entropy (MaxEnt) (Phillips, Dudík, and Schapire 2004), random forests (Valavi et al. 2021), and generalized linear models (Guisan, Edwards, and Hastie 2002). More recently, a wide variety of deep learning approaches have been introduced, including simple multi-layer perceptrons (Zbinden et al. 2024), computer vision methods leveraging remote sensing imagery (Estopinan et al. 2022; Teng et al. 2023) or location embeddings (Cole et al. 2023), and even language-based reasoning methods (Hamilton et al. 2024). Active learning (Lange et al. 2023) and few-shot learning (Lange et al. 2025) frameworks have also been introduced. Deep learning models have been shown to consistently outperform traditional approaches in ecology (Deneu et al. 2021), notably when applied to large-scale citizen science datasets (Davis et al. 2023).
Some deep learning-based SDMs have focused on estimating binary species range maps (Dorm et al. 2024). This corresponds to a multi-label classification problem in which the list of species that can be encountered at a given location is predicted, regardless of the frequency at which they are observed. Another important challenge is encounter rate estimation, i.e. a multi-output regression problem aimed at predicting the average rate at which observers encounter species at a given location (Johnston et al. 2021). The use of ML for predicting encounter rates has been explored in Teng et al. (2023) and Abdelwahed, Teng, and Rolnick (2024), a line of work which we follow here. Encounter rates are closely linked to the notion of occupancy, i.e. the probability of a species occupying a given site. Occupancy is a core concept in ecology (MacKenzie 2018), and understanding occupancy patterns is essential for prioritizing conservation efforts. Because encounter rates can be seen as the product of occupancy and detection probability (Johnston et al. 2021), they can provide a more nuanced and realistic representation of species distributions than binary range maps.
@@
In statistical ecology, SDMs have been estimated in a Bayesian context using computationally intensive algorithms, such as Markov Chain Monte Carlo (MCMC) approaches (Zulian, Miller, and Ferraz 2021) and integrated nested Laplace approximation (Omre, Fjeldstad, and Forberg 2024). Golding and Purse (2016) proposed a more flexible approach based on Gaussian processes, but this was applicable only to a limited number of species and covariates. Our work is the first to reframe deep learning-based SDMs under a Bayesian perspective.
@@
Uncertainty quantification is recognized by ecologists as a major challenge for SDMs (Yates et al. 2018). To date, however, the vast majority of works do not investigate the effects of uncertainty on model outputs (Zurell et al. 2020). Only a limited number of statistical tools have been proposed to measure and incorporate uncertainty estimates into SDMs (Beale and Lennon 2012), and these tools are not easily scalable, as they often require computationally expensive MCMC iterations to estimate posterior probability distributions (Rocchini et al. 2019). Thus, they are only appropriate for SDMs relying on a very small number of covariates and species. Approximation techniques such as model ensembling (Guo et al. 2015) have been applied to quantify uncertainty in SDMs (Convertino et al. 2012), but their use remain limited. Recently, an active learning framework was proposed to sequentially select geographic locations that would best reduce uncertainty of an SDM for previously unmapped species Lange et al. (2023). This work however does not take a Bayesian approach to improving predictions, instead estimating uncertainty via disagreement across an ensemble of models to guide the active learning step.
@@
An extensive variety of approaches have been proposed to approximate uncertainty in the Bayesian deep learning literature, since computing a posterior distribution over the parameters of a neural network according to the rule of Bayesian inference is analytically intractable (Band et al. 2021). However, these approaches have mostly been tested on generic vision benchmarks (Mucsányi, Kirchhof, and Oh 2024) such as CIFAR-10 (Krizhevsky 2009) and ImageNet-1k (Deng et al. 2009), which do not reflect the complexity of real-world tasks. Our work is the first to integrate such approaches for solving a complex ecological task.
@@ 3 The BATIS Framework
Following Teng et al. (2023); Abdelwahed, Teng, and Rolnick (2024), we consider a machine learning formulation of species distribution modeling in which the encounter rate $p^{\infty}_{i,k}$ is to be estimated for each species $i$, given a set of information associated with the hotspot $h_{k}$ – e.g., average precipitation, temperature, etc., as well as satellite imagery of the location. In practice, of course, the ground truth value $p^{\infty}_{i,k}$ is not directly observable, but is well-approximated by $p_{i,k}$ for hotspots $h_{k}$ where $|\mathcal{C}_{k}|$ is sufficiently large; therefore the task is to estimate $p_{i,k}$.
@@ 4 The BATIS Benchmark
We introduce a dataset for the BATIS framework. We rely on the eBird citizen science database (Sullivan et al. 2009), containing millions of bird observations across the globe standardized in the form of checklists, indicating which species were seen or not during a given survey trip. This large amount of data allows us to create a benchmark of test cases that mirrors the real-world scenarios in which BATIS can be applied. We can notably simulate data-deficient locations by providing partial information to SDM algorithms so as to test their generalization capabilities against full ground-truth data. Although eBird strictly focuses on birds, its structure and scale makes it an ideal setting for evaluating BATIS methods, which we hope may then be applied to less standardized datasets, such as those available through GBIF.
Following Teng et al. (2023), we first extracted complete checklists from all hotspots associated with the mainland portions of Kenya (KE), South Africa (ZA), and the contiguous United States. For the US, we followed the same approach as Teng et al. (2023), splitting the dataset into two seasons: summer (US-S, breeding period) and winter (US-W, non-breeding period). We did not consider seasonality for the KE and ZA datasets, as these regions have significantly fewer available checklists and seasonal migrations are also less pronounced in KE and ZA than in the US. Table 1 summarizes the composition of each subdataset. Further details on the dataset can be found in Appendix A.
@@
Our dataset also includes bioclimatic rasters from the WorldClim2.1 model (1km resolution) and Sentinel-2 satellite images (10m resolution) for each hotspot. These serve as predictive variables (inputs) to the models we test. More information on these variables can be found in Appendix A.
@@
We reserved 50% of the hotspots with $\geq$ 15 checklists for the test set (15 being a large enough value to reliably estimate the true encounter rate $p^{\infty}_{i,k}$). The remaining hotspots were split into training and validation sets in a 80:20 ratio, and we used DBSCAN to cluster hotspots and proportionally distribute them across splits, which prevents auto-correlation and over-fitting (Roberts et al. 2017). Table 5 (Appendix A) summarizes the content of each split.
@@
We considered four SDM baselines in our benchmark. The first one, referred to as Mean Encounter Rate, simply predicts the average encounter rate on the training set for each species, ignoring the inputs. We also include a Random Forest baseline trained from bioclimatic variables only, as this is a frequently used approach in ecology. Because bioclimatic variables are still almost always exclusively used in ecology, we include a Multi-Layer Perceptron Model (MLP) solely trained from these inputs. Finally, we consider the ResNet-18 approach introduced by Teng et al. (2023), as it was shown to be the best performing model among the SatBird benchmark. The ResNet-18 relies on both bioclimatic variables and satellite imagery as inputs. More information on each of our SDM baselines can be found in Appendix C.
@@
We considered the uncertainty estimation approaches listed below. Except for the Fixed and Historical Variance baselines, we restricted ourselves exclusively to distributional approaches, as they offer a more principled way of estimating both the mean and variance of each predicted encounter rate (Mucsányi, Kirchhof, and Oh 2024). More details on each approach can be found in Appendix C.
@@
Monte-Carlo Dropout (Gal and Ghahramani 2016).
MC Dropout (MCD) can be shown to be a variational approximation of a deep Gaussian process (Gal and Ghahramani 2016). A proportion of neurons are randomly deactivated at training and test time, and $M$ forward passes are used to compute mean and variance for each location. MCD is primarily effective at capturing epistemic uncertainty.
@@
Deep Ensembles (Lakshminarayanan, Pritzel, and Blundell 2017).
In Deep Ensembles (DE), an ensemble of $M$ neural networks are independently trained on the same task, and their individual predictions on the same input can be used to estimate mean and variance. Despite its simplicity, DE is the current state-of-the-art for estimating epistemic uncertainty in deep learning (Mukhoti et al. 2021).
Shallow-Ensembles (Lee et al. 2015).
@@
Mean-Variance Network (Nix and Weigend 1994; Sluijterman, Cator, and Heskes 2023) .
@@
Heteroscedastic Regression Neural Network (Kendall and Gal 2017).
@@
We performed the experiments described below. Training protocol details, including hyperparameters, and compute resources can be found in Appendix D.
@@
We evaluate our methods with the same metrics as Teng et al. (2023), namely MAE, MSE, Top-10, Top-30, and Top-k accuracy. Further details are provided in Appendix D.
@@ 5 Results and Discussion
Table 2 details the performance of uncertain-agnostic and uncertainty-aware models on the KE and ZA subdatasets. Results for the US can be found in Tables 10 and 11 in Appendix E. We observe that uncertainty-aware methods significantly outperforms all the uncertainty-agnostic approaches on our test set, even when prior predictions are updated with five checklists only. This suggests that our Bayesian approach to SDM estimation could greatly improve the reliability of predictions in data-scarce locations. Our proposed framework essentially merges the strengths of traditional tools for Earth-scale mapping, which can make estimates for large regions with high throughput, and of ground-based surveys, which can record specific and accurate data with high effort.
@@
Figure 2 shows the evolution of the performance of the uncertainty-aware approaches with increasing number of checklist updates, on the ZA subdataset. Similar figures for the KE and US subdatasets are provided in Appendix E. We observe that our Bayesian framework requires as few as one or two checklists to significantly outperform static baselines. MVN and HetReg are the top-performing methods in our benchmark, with HetReg performing only marginally better than MVN. This is likely because HetReg and MVN share a common aleatoric uncertainty quantification module, but HetReg also includes an epistemic uncertainty estimation module. Estimating epistemic uncertainty alone, however, does not significantly improve on static baselines, as demonstrated by the limited improvements in performance with Deep Ensembles. Our findings are consistent across all four subdatasets. Interestingly, we also observe that the Fixed Variance baseline reaches MAE and MSE performance comparable to HetReg/MVN after 10 checklists, when the prior variance is fixed to its maximum theoretical value.
@@
We find that models relying on aleatoric uncertainty (such as HetReg and MVN) perform better than those which rely on epistemic uncertainty (such as DE and MCD). This likely results from two factors: First, with real-world observation-based data, similar conditions can occur with very different observations (in this case, reflecting variability in the whereabouts of individual animals and their visibility to an observer). Second, subtle environmental differences may not be fully captured by the input variables. Our conclusions are consistent with trends noted in the Bayesian ML literature (Kendall and Gal 2017), in which it is observed that epistemic uncertainty can be poorly estimated in deep learning models due to the implicit assumption that similar inputs lead to similar outcomes. By contrast, aleatoric uncertainty estimation approaches allow us to capture the inherent unpredictability of species observations and the potential existence of local environmental factors not captured by the prior SDM.

### v2

@@ Abstract
Codehttps://github.com/cath34/batis
Datasetshttps://huggingface.co/datasets/cathv/BATIS
Extended versionhttps://arxiv.org/abs/2510.19749
@@ 1 Introduction
Understanding the factors governing the spatial distribution of species is a long-standing topic of interest in ecology (decandolle; pollock2015). The rapid loss of biodiversity further highlights the need to better understand species distributions and conservation status, which requires robust modeling methods (nospeciesleftbehind). Species distribution models (SDMs) are at the forefront of this research, enabling ecologists to predict and understand species occurrence by linking observational data (presence or absence) to environmental covariates (elithsdm). These models have proved to be essential tools to support conservation planning and management (sdmessentialtool), with notable examples including predicting species range shift (rangeshiftpaper), mapping invasive species risk (alienspecies), supporting translocation programs (translocpaper), reducing bycatch rates in fisheries (fisheriespaper), and informing IUCN Red List assessments (iucnpaper).
SDMs are traditionally based on statistical approaches, offering strong interpretability and theoretical grounding. However these methods often rely on restrictive assumptions such as linearity, leading them to struggle with high-dimensional datasets. Recently, deep learning approaches to SDMs proposed to overcome these challenges and were shown to consistently outperform statistical ecology methods for large and heterogeneous datasets (Deneu_Servajean_Bonnet_Botella_Munoz_Joly_2021).
Although a wide variety of SDM approaches have now been introduced, their effectiveness at large spatio-temporal scales remains hindered by the abundance and geographical coverage of the data used to train them (minimumsdmdata). In most cases, observations are disproportionately concentrated in areas easily accessible by roads (roadbias) and with high human population densities (densitybias). In contrast, remote locations are most often associated with very limited to no observations, as illustrated in Fig. 1.
@@
Data distribution biases especially affect machine learning-based approaches, which typically rely on trustworthy labeled data. Limited observations in a given location can lead to poor estimates of species occurrence (finkpaperencounterate) – essentially, noisy labels – and naïvely incorporating such data points in model training can degrade performance. However, sparse observations still offer potentially valuable information, and finding ways to properly leverage it could help improve SDMs and guide conservation efforts in remote or underresourced areas. We argue that uncertainty-aware methods can offer a principled way of leveraging limited location-specific observations, enabling controlled prediction updates that can balance global-scale models with local-scale observations.
@@
  • We benchmark a variety of state-of-the-art uncertainty estimation ML methods on the above framework, and present a novel dataset derived from eBird (ebirdpaper), building on the SatBird (NEURIPS2023_ef7653bb) dataset.
@@ 2 Related Works
Species distributions have been widely estimated from environmental data with statistical ecology methods. Popular SDMs in ecology include techniques based on maximum entropy (MaxEnt) (maxenticml), random forests (rfsdm), and generalized linear models (GUISAN200289). More recently, a wide variety of deep learning approaches have been introduced, including simple multi-layer perceptrons (zbindenpresenceabsence), computer vision methods leveraging remote sensing imagery (Estopinan_Servajean_Bonnet_Munoz_Joly_2022; NEURIPS2023_ef7653bb) or location embeddings (SINR_icml23), and even language-based reasoning methods (NEURIPS2024_1f96b24d). Active learning (NEURIPS2023_82eec786) and few-shot learning (fewshotpaper) frameworks have also been introduced. Deep learning models have been shown to consistently outperform traditional approaches in ecology (Deneu_Servajean_Bonnet_Botella_Munoz_Joly_2021), notably when applied to large-scale citizen science datasets (gomescitizen).
Some deep learning-based SDMs have focused on estimating binary species range maps (dorm2024generatingbinaryspeciesrange). This corresponds to a multi-label classification problem in which the list of species that can be encountered at a given location is predicted, regardless of the frequency at which they are observed. Another important challenge is encounter rate estimation, i.e. a multi-output regression problem aimed at predicting the average rate at which observers encounter species at a given location (finkebird1). The use of ML for predicting encounter rates has been explored in NEURIPS2023_ef7653bb and abdelwahed2024predictingspeciesoccurrencepatterns, a line of work which we follow here. Encounter rates are closely linked to the notion of occupancy, i.e. the probability of a species occupying a given site. Occupancy is a core concept in ecology (MacKenzie_2018), and understanding occupancy patterns is essential for prioritizing conservation efforts. Because encounter rates can be seen as the product of occupancy and detection probability (finkebird1), they can provide a more nuanced and realistic representation of species distributions than binary range maps.
@@
In statistical ecology, SDMs have been estimated in a Bayesian context using computationally intensive algorithms, such as Markov Chain Monte Carlo (MCMC) approaches (Zulian_Miller_Ferraz_2021) and integrated nested Laplace approximation (Omre_Fjeldstad_Forberg_2024). Golding_Purse_2016 proposed a more flexible approach based on Gaussian processes, but this was applicable only to a limited number of species and covariates. Our work is the first to reframe deep learning-based SDMs under a Bayesian perspective.
@@
Uncertainty quantification is recognized by ecologists as a major challenge for SDMs (importantuncertaintysdm). To date, however, the vast majority of works do not investigate the effects of uncertainty on model outputs (paperodmap). Only a limited number of statistical tools have been proposed to measure and incorporate uncertainty estimates into SDMs (uncsdm1), and these tools are not easily scalable, as they often require computationally expensive MCMC iterations to estimate posterior probability distributions (Rocchini_Marcantonio_Arhonditsis_Cacciato_Hauffe_He_2019). Thus, they are only appropriate for SDMs relying on a very small number of covariates and species. Approximation techniques such as model ensembling (GUO201567) have been applied to quantify uncertainty in SDMs (CONVERTINO20121), but their use remain limited. Recently, an active learning framework was proposed to sequentially select geographic locations that would best reduce uncertainty of an SDM for previously unmapped species NEURIPS2023_82eec786. This work however does not take a Bayesian approach to improving predictions, instead estimating uncertainty via disagreement across an ensemble of models to guide the active learning step.
@@
An extensive variety of approaches have been proposed to approximate uncertainty in the Bayesian deep learning literature, since computing a posterior distribution over the parameters of a neural network according to the rule of Bayesian inference is analytically intractable (benchmarkdiabetes). However, these approaches have mostly been tested on generic vision benchmarks (NEURIPS2024_5afa9cb1) such as CIFAR-10 (Krizhevsky2009LearningML) and ImageNet-1k (5206848), which do not reflect the complexity of real-world tasks. Our work is the first to integrate such approaches for solving a complex ecological task.
@@ 3 The BATIS Framework
Following NEURIPS2023_ef7653bb; abdelwahed2024predictingspeciesoccurrencepatterns, we consider a machine learning formulation of species distribution modeling in which the encounter rate $p^{\infty}_{i,k}$ is to be estimated for each species $i$, given a set of information associated with the hotspot $h_{k}$ – e.g., average precipitation, temperature, etc., as well as satellite imagery of the location. In practice, of course, the ground truth value $p^{\infty}_{i,k}$ is not directly observable, but is well-approximated by $p_{i,k}$ for hotspots $h_{k}$ where $|\mathcal{C}_{k}|$ is sufficiently large; therefore the task is to estimate $p_{i,k}$.
@@ 4 The BATIS Benchmark
We introduce a dataset for the BATIS framework. We rely on the eBird citizen science database (ebirdpaper), containing millions of bird observations across the globe standardized in the form of checklists, indicating which species were seen or not during a given survey trip. This large amount of data allows us to create a benchmark of test cases that mirrors the real-world scenarios in which BATIS can be applied. We can notably simulate data-deficient locations by providing partial information to SDM algorithms so as to test their generalization capabilities against full ground-truth data. Although eBird strictly focuses on birds, its structure and scale makes it an ideal setting for evaluating BATIS methods, which we hope may then be applied to less standardized datasets, such as those available through GBIF.
Following NEURIPS2023_ef7653bb, we first extracted complete checklists from all hotspots associated with the mainland portions of Kenya (KE), South Africa (ZA), and the contiguous United States. For the US, we followed the same approach as NEURIPS2023_ef7653bb, splitting the dataset into two seasons: summer (US-S, breeding period) and winter (US-W, non-breeding period). We did not consider seasonality for the KE and ZA datasets, as these regions have significantly fewer available checklists and seasonal migrations are also less pronounced in KE and ZA than in the US. Table 1 summarizes the composition of each subdataset. Further details on the dataset can be found in Appendix A.
@@
Our dataset also includes bioclimatic rasters from the WorldClim2.1 model (1km resolution) and Sentinel-2 satellite images (10m resolution) for each hotspot. These serve as predictive variables (inputs) to the models we test, and were selected based on the work of NEURIPS2023_ef7653bb. More information on these variables can be found in Appendix A.
@@
We reserved 50% of the hotspots with $\geq$ 15 checklists for the test set (15 being a large enough value to reliably estimate the true encounter rate $p^{\infty}_{i,k}$). The remaining hotspots were split into training and validation sets in a 80:20 ratio, and we used DBSCAN to cluster hotspots and proportionally distribute them across splits, which prevents auto-correlation and over-fitting (autocorrgis). Table 5 (Appendix A) summarizes the content of each split.
@@
We considered four SDM baselines in our benchmark. The first one, referred to as Mean Encounter Rate, simply predicts the average encounter rate on the training set for each species, ignoring the inputs. We also include a Random Forest baseline trained from bioclimatic variables only, as this is a frequently used approach in ecology. Because bioclimatic variables are still almost always exclusively used in ecology, we include a Multi-Layer Perceptron Model (MLP) solely trained from these inputs. Finally, we consider the ResNet-18 approach introduced by NEURIPS2023_ef7653bb, as it was shown to be the best performing model among the SatBird benchmark. The ResNet-18 relies on both bioclimatic variables and satellite imagery as inputs. More information on each of our SDM baselines can be found in Appendix F.
@@
We considered the uncertainty estimation approaches listed below. Except for the Fixed and Historical Variance baselines, we restricted ourselves exclusively to distributional approaches, as they offer a more principled way of estimating both the mean and variance of each predicted encounter rate (NEURIPS2024_5afa9cb1). More details on each approach can be found in Appendix F.
@@
Monte-Carlo Dropout (pmlr-v48-gal16).
MC Dropout (MCD) can be shown to be a variational approximation of a deep Gaussian process (pmlr-v48-gal16). A proportion of neurons are randomly deactivated at training and test time, and $M$ forward passes are used to compute mean and variance for each location. MCD is primarily effective at capturing epistemic uncertainty.
@@
Deep Ensembles (deepens).
In Deep Ensembles (DE), an ensemble of $M$ neural networks are independently trained on the same task, and their individual predictions on the same input can be used to estimate mean and variance. Despite its simplicity, DE is the current state-of-the-art for estimating epistemic uncertainty in deep learning (mukhoti2021deterministic).
Shallow-Ensembles (lee2015mheadsbetterone).
@@
Mean-Variance Network (mvapaper2; mvapaper1) .
@@
Heteroscedastic Regression Neural Network (hetregpaper).
@@
We performed the experiments described below. Training protocol details, including hyperparameters, and compute resources can be found in Appendix G.
@@
We evaluate our methods with the same metrics as NEURIPS2023_ef7653bb, namely MAE, MSE, Top-10, Top-30, and Top-k accuracy. Further details on these metrics are provided in Appendix G.
@@ 5 Results and Discussion
Table 2 details the performance of uncertain-agnostic and uncertainty-aware models on the KE and ZA subdatasets. Results for the US can be found in Tables 10 and 11 in Appendix H. We observe that uncertainty-aware methods significantly outperforms all the uncertainty-agnostic approaches on our test set, even when prior predictions are updated with five checklists only. This suggests that our Bayesian approach to SDM estimation could greatly improve the reliability of predictions in data-scarce locations. Our proposed framework essentially merges the strengths of traditional tools for Earth-scale mapping, which can make estimates for large regions with high throughput, and of ground-based surveys, which can record specific and accurate data with high effort.
@@
Figure 2 shows the evolution of the performance of the uncertainty-aware approaches with increasing number of checklist updates, on the ZA subdataset. Similar figures for the KE and US subdatasets are provided in Appendix H. We observe that our Bayesian framework requires as few as one or two checklists to significantly outperform static baselines. MVN and HetReg are the top-performing methods in our benchmark, with HetReg performing only marginally better than MVN. This is likely because HetReg and MVN share a common aleatoric uncertainty quantification module, but HetReg also includes an epistemic uncertainty estimation module. Estimating epistemic uncertainty alone, however, does not significantly improve on static baselines, as demonstrated by the limited improvements in performance with Deep Ensembles. Our findings are consistent across all four subdatasets. Interestingly, we also observe that the Fixed Variance baseline reaches MAE and MSE performance comparable to HetReg/MVN after 10 checklists, when the prior variance is fixed to its maximum theoretical value.
@@
We find that models relying on aleatoric uncertainty (such as HetReg and MVN) perform better than those which rely on epistemic uncertainty (such as DE and MCD). This likely results from two factors: First, with real-world observation-based data, similar conditions can occur with very different observations (in this case, reflecting variability in the whereabouts of individual animals and their visibility to an observer). Second, subtle environmental differences may not be fully captured by the input variables. Our conclusions are consistent with trends noted in the Bayesian ML literature (hetregpaper), in which it is observed that epistemic uncertainty can be poorly estimated in deep learning models due to the implicit assumption that similar inputs lead to similar outcomes. By contrast, aleatoric uncertainty estimation approaches allow us to capture the inherent unpredictability of species observations and the potential existence of local environmental factors not captured by the prior SDM.
@@
6 Limitations
There remains tension between the broad environmental variables learned by deep SDMs (our Bayesian prior) and the hotspot-specific features captured by observational information (our updates to that prior). As shown in Figure 2, the initial predictions of an uncertainty-aware SDM can be overwritten in a first Bayesian update, when the estimated prior variance is large. This can lead to an initial drop in performance before enough updates are performed to boost performance, even though performance would ideally increase monotonically with number of updates. A simple approach to mitigate such behavior is to consider a weighted average of predictions from the prior and the updated SDM, progressively giving more importance to the updates with more observations. We provide additional experiments showing how such an approach can help smooth the results curve observed in Fig. 2 in Appendix E..
Citizen science data is heavily influenced by predictable factors such as the day of the week, the time of day, the expertise of the observer, the weather, and the proximity to population centers (sierra2025divshift). Such biases are structured, which technically violates the randomness assumption of aleatoric uncertainty. Epistemic uncertainty estimation approaches can potentially correct for biases if they can be explicitly modeled, but are unable to do so if they are not encoded in the data. While our work shows that is possible to compensate partially for data gaps associated with the irregular distribution of citizen science data, some biases may affect the accuracy of the underlying numbers used as ground truth, which cannot be compensated for by such methods and remain an active object of inquiry in statistical ecology (spatialbiasref).